OpenStudy (anonymous):

8x+30=4x+18 what is x

OpenStudy (anonymous):

\[8x+30=4x+18\] \[8x-4x=18-30\] \[4x=-12\] \[x=\frac{-12}{4}=-3\]

OpenStudy (anonymous):

the answer is 11

OpenStudy (anonymous):

\[8\times 11+30=4\times 11+18\] \[88+30=44+18\] \[118=62\] no 11 does not work

OpenStudy (anonymous):

\[8\times -3+30=4\times -3+18\] \[-24+30=-12+18\] \[6=6\] i am sticking with \[x=-3\]

OpenStudy (anonymous):

i dont know how but when i typed in that answer it said the correct answer was 11

OpenStudy (anonymous):

sure you copied down the question correctly?

OpenStudy (anonymous):

yes its a parallelogram

OpenStudy (anonymous):

what is x when 116 + 4x+32